OpenStudy (anonymous):

my answer is C Solve 4x + 7 = 6x - 1 (2 points) Answer x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (anonymous):

@phi

OpenStudy (johnweldon1993):

is this written as \[4^{x+7} = 6^{x-1} ?\]

OpenStudy (anonymous):

yes @johnweldon1993

OpenStudy (phi):

can you re-write it using the rule \[ log(a^{compicated\ stuff})= (compicated\ stuff) \log(a)\]

OpenStudy (johnweldon1993):

lol @phi nice use of "complicated stuff"

OpenStudy (phi):

in other words, take the log of both sides: \[ \log({4^{x+7}}) =\log{( 6^{x-1}}) \]

OpenStudy (phi):

now use the rule. what do you get ?

OpenStudy (anonymous):

hahahahahahahaha complicated stuff

OpenStudy (anonymous):

My answer now is D

OpenStudy (phi):

how did you get it ?

OpenStudy (anonymous):

I solved for x

OpenStudy (anonymous):

Carrying all numbers to 5 significant figures,